From a height of 1.5 m above the surface of the moon (no friction) a penny is dropped. The travel time and the velocity at arrival on the ground are ...?
Answer #1

Acceleration due to gravity at moon, a = 1.622 m/s^2
h = 1.5 m

h = u*t + 1/2*a*t^2
1.5 = 0 + 1/2*1.622*t^2
t = 1.36 s

V = u + a*t
V = 0 + 1.622 * 1.36
V = 2.21 m/s